OpenStudy (anonymous):

\(\large p\) is a prime number such that there exist positive integers \(\large n\) and \(\large m\) that satisfy the following equation: \(\Large p^n+144=m^2\) What is the sum of all possible values of \(\large m\) ?

OpenStudy (anonymous):

\(\huge Solution\) Rewriting gives: \(\Large p^n=(m+12)(m-12)\). Let \(\Large m+12=p^i\) and \(\Large m-12=p^j\). So \(\Large i+j=n\) and \(\Large i\gt j\) Then we have: \(\Large p^j(p^{i-j}-1)=p^i-p^j=(m+12)-(m-12)=24\\\Large =5^0(5^2-1)=2^3(2^2-1)=3^1(3^2-1)\) From here we have 3 solutions: \(\Large m=13,15,20\) with sum \(\huge 48\)

OpenStudy (vishweshshrimali5):

Great work @JungHyunRan :)

OpenStudy (anonymous):

thanks :)

ganeshie8 (ganeshie8):

brilliant ! xD